Solve for the determinant in the equation below(10marks)

1.7.1

"\begin{matrix}
 4 & -3 & 2 \\
 1 & 2 &-2 \\
 2 & -1 & -4\\

\end{matrix}"

1.7.2.

"\begin{matrix}
 2 & -2 & 1 \\
 2 & 2 & 1\\
 4 & 1 & 3 \\
\end{matrix}"


1
Expert's answer
2022-04-18T00:25:59-0400

1.7.1.

"det A =4\u00b72\u00b7(-4) + (-3)\u00b7(-2)\u00b72 + 2\u00b71\u00b7(-1) - 2\u00b72\u00b72 - 4\u00b7(-2)\u00b7(-1) - (-3)\u00b71\u00b7(-4) = -32 + 12 - 2 - 8 -\n\n- 8 - 12 = -50"


1.7.2.

"det A = 2\u00b72\u00b73 + (-2)\u00b71\u00b74 + 1\u00b72\u00b71 - 1\u00b72\u00b74 - 2\u00b71\u00b71 - (-2)\u00b72\u00b73 = 12 - 8 + 2 - 8 - 2 + 12 = 8"